solve by cramer's rule:1/x+1/y+1/z=1,2/x+5/y+3/z=0,1/x+2/y+4/z=3

Answer by stanbon(75887)

solve by cramer's rule:
1/x+1/y+1/z=1
2/x+5/y+3/z=0
1/x+2/y+4/z=3
-----------
I used a Matrix function on my TI-84 to get:
1/x = 1
1/y = -1
1/z = 1
===========
Cheers,
Stan H.
===========